# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 6. The Invariants of ${{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}$
Using Lemma 6.1, $u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}\in R_{1}$. Since $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}$ and $u_{2}d_{2,2}-\xi_{1}^{q}\xi_{3}$ are elements of $R_{2}$,
$u_{2}\xi_{3}-(\xi_{2}+c_{2,2})u_{2}d_{1,2}+\xi_{1}u_{2}d_{2,2}\in R_{2}.$
Since this expression is divisible by $u_{2}$ in $S_{2}$ and lies in $R_{2}$, it is divisible by $u_{2}$ in $R_{2}$. Defining $-c_{3,2}:=(u_{2}\xi_{3}-(\xi_{2}+c_{2,2})u_{2}d_{1,2}-\xi_{1}u_{2}d_{2,2})/u_{2}$ gives $\xi_{3}=(\xi_{2}+c_{2,2})d_{1,2}-\xi_{1}d_{2,2}-c_{3,2}$ with $c_{3,2}\in R_{2}$. Using Lemma 6.4
$(\xi_{2}+c_{2,2})u_{2}d_{1,2}-\xi_{1}u_{2}d_{2,2}$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{3}((\xi_{2}+c_{2,2})\xi_{0}^{q}-\xi_{1}^{q+1})-\xi_{2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$+\xi_{1}(\xi_{2}^{q+1}+\xi_{0}^{q}\xi_{1}^{q^{2}-q}\xi_{2})-c_{2,2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$=$ $u_{2}\xi_{3}+\xi_{0}\xi_{1}^{q}\xi_{2}^{q}+\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}$
$-c_{2,2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
Therefore, using Lemma 6.1 to substitute for $c_{2,2}$,
$-u_{2}c_{3,2}$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $-\xi_{0}\xi_{1}^{q}\xi_{2}^{q}-\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}+c_{2,2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$\equiv_{\langle\xi_{0}^{q+2}\rangle}$ $-\xi_{0}\xi_{1}^{q}\xi_{2}^{q}-\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}+\xi_{0}\xi_{1}^{q-1}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$=$ $-\xi_{0}^{2}\xi_{1}^{2q-1}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}$
$=$ $-\xi_{1}^{q+1}(\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}).$
Thus $-c_{3,2}\equiv_{\langle\xi_{0}^{q+2}\rangle}\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}.$
Applying $\mathcal{P}^{1}$ to $\xi_{3}=(\xi_{2}+c_{2,2})d_{1,2}-\xi_{1}d_{2,2}-c_{3,2}$ and using $\mathcal{P}^{1}(c_{2,2})=\xi_{1}^{q}$ from Lemma 6.3 gives
$\xi_{2}^{q}=2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}-\mathcal{P}^{1}(c_{3,2})+(\xi_{2}+c_{2,2})\mathcal{P}^{1}(d_{1,2})-\xi_{1}\mathcal{P}^{1}(d_{2,2}).$
Since $c_{3,2}\in R_{2}$, $\mathcal{P}^{1}(c_{3,2})\in R_{2}$. Furthermore, using Lemma 6.8,
$\mathcal{P}^{1}(c_{3,2})\equiv_{\langle\xi_{0}^{q+2}\rangle}\mathcal{P}^{1}(\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2})=2\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}+\xi_{0}^{q}\xi_{1}^{q^{2}-q}$
Using Lemmas 6.4 and 6.5, $\mathcal{P}^{1}(u_{2}d_{1,2})=u_{2}\mathcal{P}^{1}(d_{1,2})$, $\mathcal{P}^{1}(d_{1,2})\in R_{2}$ and
$u_{2}\mathcal{P}^{1}(d_{1,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{0}^{q}\xi_{2}^{q}-2\xi_{0}^{q}\xi_{2}^{q}+\xi_{1}^{q+1}\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{2q}\xi_{2}^{q-2}=\xi_{1}^{q+1}\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{2q}\xi_{2}^{q-2}-\xi_{0}^{q}\xi_{2}^{q}.$
Again using Lemmas 6.4 and 6.5, $\mathcal{P}^{1}(u_{2}d_{2,2})=u_{2}\mathcal{P}^{1}(d_{1,2})$, $\mathcal{P}^{1}(d_{2,2})\in R_{2}$ and $u_{2}\mathcal{P}^{1}(d_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{2}^{q}-\xi_{1}\xi_{2}^{q}-\xi_{0}^{q}\xi_{1}^{q^{2}}=-\xi_{0}^{q}\xi_{1}^{q^{2}}$. Since $u_{2}\equiv_{\langle\xi_{0}^{q+1}\rangle}-\xi_{1}^{q+1}+\xi_{0}^{q}\xi_{2}$, we have
$\mathcal{P}^{1}(d_{1,2})\equiv_{\langle\xi_{0}^{q+1}\rangle}-\xi_{2}^{q-1}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2}$
and $\mathcal{P}^{1}(d_{2,2})\equiv_{\langle\xi_{0}^{q+1}\rangle}\xi_{0}^{q}\xi_{1}^{q^{2}-q-1}$. Since $\mathcal{P}^{1}(d_{1,2})$ and $\mathcal{P}^{1}(d_{2,2})$ lie in $R_{2}$, we have $\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}\in R_{2}$. Using the congruences we have
$\xi_{2}^{q}$ $\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}+(2\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}-\xi_{0}^{q}\xi_{1}^{q^{2}-q})$
$-(\xi_{2}+c_{2,2})(\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2})+\xi_{0}^{q}\xi_{1}^{q^{2}-q}$
$\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}+3\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{2}^{q}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}$
$-c_{2,2}(\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2})$
$\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}-\xi_{2}^{q}+3\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}$
$-\xi_{0}\xi_{1}^{q-1}(\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2})$
$\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}-\xi_{2}^{q}+2\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}.$
@@
$\xi_{2}^{q}\equiv_{\langle\xi_{0}^{q+1}\rangle}\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}.$
Recall that $-\xi_{1}^{q}d_{1,2}+\xi_{0}^{q}d_{2,2}\in R_{2}$ and define $r_{2,1}\in R_{2}$ by
(3) $\xi_{0}^{q+1}r_{2,1}=\xi_{2}^{q}-\xi_{1}^{q}d_{1,2}+\xi_{0}^{q}d_{2,2}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}.$
Then
(4) $\xi_{2}^{q}=\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}-\gamma_{2,2}^{(1)}$
with $\gamma_{2,2}^{(1)}=-\xi_{0}^{q+1}r_{2,1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}\in R_{2}$ and $\nu(\gamma_{2,2}^{(1)})>q+1$. Thus part (e) of Theorem 4.6 is satisfied for $m=2$.
Lemma 7.2 does not rely on induction and the proof is valid for $m=2$. This completes the proof of Theorem 4.6 for $m=2$.
Remark 6.9.
Notice that in the process of deriving Equation (4), the expression $-\xi_{2}^{q}$ arises on the right hand side. If the coefficent had been +1 rather than -1 then this process would not have provided a formula for $\xi_{2}^{q}$. It is for this reason that naively applying a Steenrod operation to the formula for $\xi_{n-1-i}^{q^{i}}$ is not guaranteed to give a formula for $\xi_{n-2-i}^{q^{i+1}}$.
7. Proof of the Main Theorem
The proof is by induction on $m$. The base case, $m=2$, is given in Section 6. For the induction step, we will assume the theorem is true for $S_{\ell}^{G_{\ell}}$ for $2\leq\ell<m$ and, in the proof of a given part of the theorem, we will assume the earlier parts are true for $S_{m}^{G_{m}}$.
Remark 7.1.
Part (c) is Theorem 3.11.
Lemma 7.2.
Part (f) of Theorem 4.6 follows from Parts (a)–(e).
Proof.
Let $A$ denote the subalgebra of $S_{m}^{G_{m}}$ generated by $\mathcal{H}\cup\{\xi_{m},\ldots,\xi_{n-2}\}$. Using Part (c), $A\subseteq S_{m}^{G_{m}}$ is an integral extension. Using Part (d), $R_{n-1}\subset A$. Therefore $\mathcal{F}(A)=\mathcal{F}(R_{n-1})=\mathcal{F}(S_{m}^{G_{m}})$. Thus to show $A=S_{m}^{G_{m}}$, it is sufficient to show that $A$ is integrally closed in its field of fractions. We will use [17, Proposition 1.1] to prove that $A$ is a UFD. Recall that a UFD is integrally closed in its field of fractions (see, for example, [13, Proposition 4.10]).
Using the definition and part (e), $A$ is the $\mathbb{F}_{q}[\mathcal{H}]$-module generated by $\mathcal{B}$. Note that the number of elements of the set $\mathcal{B}$ is $\prod_{i=1}^{m-1}q^{i}=q^{m(m-1)/2}$. We know that $S_{m}^{G_{m}}$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module of rank
which simplifies to $q^{m(m-1)/2}$. Therefore the degree of the field extension $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])\subset\mathcal{F}(S_{m}^{G_{m}})$ is $q^{m(m-1)/2}$. Since $\mathcal{F}(A)=\mathcal{F}(S_{m}^{G_{m}})$, the degree of the field extension $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])\subset\mathcal{F}(A)$ is $q^{m(m-1)/2}$. We know that $\mathcal{B}$ is a spanning set for $\mathcal{F}(A)$ over $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])$. By comparing the order of $\mathcal{B}$ with the degree of the extension, we see that $\mathcal{B}$ is linearly independent over $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])$. Therefore $A$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$ and $A$ is Cohen-Macaulay. By comparing Hilbert series, we see that $A$ is a complete intersection with relations given by Part (e). Therefore, to complete the proof of Part (f), we need only show that $A=S_{m}^{G_{m}}$.
From Part (b) for $m$ and $m-1$, we have $\operatorname{LT}(u_{m})=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1}+q^{m-2}+\cdots+1}$ and $\operatorname{LT}(u_{m-1})=(-1)^{\lfloor(m-1)/2\rfloor}\xi_{m-2}^{q^{m-2}+q^{m-3}+\cdots+1}$. Therefore, since $\mathcal{H}$ is an hsop (see Part (c)), $\{u_{m-1},u_{m}\}$ is a partial hsop. Since $A$ is Cohen-Macaulay, this means that $[u_{m-1},u_{m}]$ is a regular sequence in $A$. From (a) for $m-1$, we have $\ker(\Phi_{n-4,m-2})=u_{m-1}R_{n-4}$. Therefore $u_{m-1}$ is prime in $R_{n-4}$. Since $R_{n-1}$ is a polynomial extension of $R_{n-4}$, $u_{m-1}$ is prime in $R_{n-1}$ and in its localisation $R_{n-1}[u_{m}^{-1}]$. Using Part (d), $R_{n-1}\subset A$ and by definition $u_{m}d_{i,m}\in R_{n-1}$. Therefore $A[u_{m}^{-1}]=R_{n-1}[u_{m}^{-1}]$ and $u_{m-1}$ is prime in $A[u_{m}^{-1}]$. To satisfy the hypotheses of [17, Proposition 1.1] and complete the proof, we need to show that $A[u_{m-1}^{-1}]$ is a UFD.
From Part (a), we have $u_{m}d_{1,m}-u_{m-1}^{q}\xi_{n-1}\in R_{n-2}$. Therefore $R_{n-1}\subset R_{n-2}[d_{1,m},u_{m-1}^{-1}]$ and $\mathcal{F}(R_{n-1})=\mathcal{F}(R_{n-2}[d_{1,m}])$. Hence $R_{n-2}[d_{1,m}]$ is a polynomial algebra and $R_{n-2}[d_{1,m},u_{m-1}^{-1}]$ is a UFD. We prove $A[u_{m-1}^{-1}]$ is a UFD by showing $A[u_{m-1}^{-1}]=R_{n-2}[d_{1,m},u_{m-1}^{-1}]$. By definition, $R_{n-2}[d_{1,m}]\subset A$. From Part (a), we have $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$ for $i>1$. Cross-multiplying to eliminate $\xi_{n-1}$ gives
$u_{m-1}^{q}u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}u_{m}d_{1,m}\in R_{n-2}.$
Using $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$ (from(a)), we see that
$u_{m-1}^{q}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}d_{1,m}\in R_{n-2}$
(with $u_{m-1}d_{i-1,m-1}\in R_{n-3}$). Thus $d_{i,m}\in R_{n-2}[d_{1,m},u_{m-1}^{-1}]$, as required. ∎
Lemma 7.3.
Part (d) of Theorem 4.6 follows from Part (a).
Proof.
From (a) we have
$u_{m}=(\xi_{n-2}+c_{n-2,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(u_{m-1}d_{i,m-1})^{q}$
with $c_{j,m}\in R_{j-1}$ and $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$. Therefore
$\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-1-i}+c_{n-1-i,m})(u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1})\in R_{n-2}.$
Note that the coefficient of $\xi_{n-1}$ in this expression is $u_{m}$. Dividing by $u_{m}$ gives
$c_{n-1,m}:=\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-1-i}+c_{n-1-i,m})d_{i,m}\right)-\xi_{n-1}\in R_{n-2}$
as required. ∎
7.4. Proof of Part (a)
The proof is by induction on $m$. The base cases $m=2$ is given in Section 6. By the induction hypothesis $S_{m-1}^{G_{m-1}}$ is generated by $\{\xi_{0},\ldots,\xi_{n-4},d_{1,m-1},\ldots,d_{m-1,m-1}\}$. The Hook subgroup for $G_{m}$ is a normal subgroup of the pointwise stabiliser $(G_{m})_{x_{1}}$. If we identify $G_{m-1}$ with the pointwise stabiliser of $\{x_{1},y_{1}\}$ in $G_{m}$, then $(G_{m})_{x_{1}}=G_{m-1}H$. Arguing as in the proof of Theorem 5.9 and using Theorem 5.4, we see that $S_{m}^{(G_{m})_{x_{1}}}$ is generated by
$\{N(y_{1}),x_{1},\xi_{0},\ldots,\xi_{n-3},\psi_{1}(\sigma(d_{1,m-1})),\ldots\psi_{1}(\sigma(d_{m-1,m-1}))\}.$
In this context, the compliance relation (Lemma 5.2) allows us to rewrite $\xi_{n-2}$ in terms of the generators of $S_{m}^{(G_{m})_{x_{1}}}$. For degree reasons, the resulting expression is linear in the $\psi_{1}(\sigma(d_{j,m-1}))$, giving
$\xi_{n-2}=x_{1}N(y_{1})+\lambda_{0}+\sum_{j=1}^{m-1}\lambda_{j}\psi_{1}(\sigma(d_{j,m-1}))$
with $\lambda_{j}\in R_{n-3}[x_{1}]$. Since $u_{m}=x_{1}N(y_{1})\psi_{1}(\sigma(u_{m-1}))$, we have
$u_{m}=\psi_{1}(\sigma(u_{m-1}))\left(\xi_{n-2}-\lambda_{0}-\sum_{j=1}^{m-1}\lambda_{j}\psi_{1}(\sigma(d_{j,m-1}))\right).$
We know that $\psi_{1}$ and $\sigma$ are algebra maps. Furthermore, $\psi_{1}(\sigma(f))=\psi_{1}(f)$ for $f\in R_{n-3}$. By induction $u_{m-1}$ and $u_{m-1}d_{j,m-1}$lie in $R_{n-3}$. Therefore
$u_{m}=(\xi_{n-2}-\lambda_{0})\psi_{1}(u_{m-1})-\sum_{j=1}^{m-1}\lambda_{j}\psi_{1}(u_{m-1}d_{j,m-1}).$
Write $\lambda_{j}=\overline{\lambda_{j}}+x_{1}\lambda_{j}^{\prime}$ with $\overline{\lambda_{j}}\in R_{n-3}$ and $\lambda_{j}^{\prime}\in R_{n-3}[x_{1}]$ and observe that $\psi_{1}(\lambda_{j})=\psi_{1}(\overline{\lambda_{j}})$. Using Lemma 4.2, $\psi_{1}(\psi_{1}(f))=\psi_{1}(f)^{q}$ for $f\in R_{n-3}$. Applying $\psi_{1}$ to the above expression and using the fact that $\psi_{1}(u_{m})=0$ gives
$0=\psi_{1}(\xi_{n-2}-\overline{\lambda_{0}})\psi_{1}(u_{m-1})^{q}-\sum_{j=1}^{m-1}\psi_{1}(\overline{\lambda_{j}})\psi_{1}(u_{m-1}d_{j,m-1})^{q}.$
Therefore
$(\xi_{n-2}-\overline{\lambda_{0}})u_{m-1}^{q}-\sum_{j=1}^{m-1}\overline{\lambda_{j}}(u_{m-1}d_{j,m-1})^{q}\in R_{n-2}\cap\ker(\psi_{1}).$
Using Lemma 4.2, $u_{m}$ divides every element of $R_{n-2}\cap\ker(\psi_{1})$. By comparing lead terms, we conclude
$u_{m}=(\xi_{n-2}-\overline{\lambda_{0}})u_{m-1}^{q}-\sum_{j=1}^{m-1}\overline{\lambda_{j}}(u_{m-1}d_{j,m-1})^{q}\in R_{n-2}$
and $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$. Applying $\mathcal{P}^{e(i,m)}$ to this expression shows that $u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(u_{m})\in R_{n-1}$ and $u_{m}d_{i,m}-\xi_{n-1}\mathcal{P}^{e(i,m)-q^{n-2}}(u_{m-1}^{q})\in R_{n-2}$. Since $e(i,m)-q^{n-2}=qe(i-1,m-1)$, we have $u_{m}d_{i,m}-\xi_{n-1}(u_{m-1}d_{i-1,m-1})^{q}\in R_{n-2}$.
@@ 7.4. Proof of Part (a)
As an element of $S_{m}$, $u_{m}$ is a product of linear forms. Therefore $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $S_{m}$. If $\mathcal{P}^{i}(u_{m})\in R_{n-2}$, then $\mathcal{P}^{i}(u_{m})\in\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$ and $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $R_{n-2}$. For $i<q^{n-2}$, $\mathcal{P}^{i}(u_{m})\in R_{n-2}$ and, therefore, $\mathcal{P}^{i}(u_{m})/u_{m}\in R_{n-2}$ By induction, $\mathcal{P}^{i}(u_{m-1})=0$ for $0<i<q^{m-2}$. Therefore, for $0<i<q^{m-1}$, $\mathcal{P}^{i}(u_{m-1}^{q})=0$ and $\mathcal{P}^{i}(u_{m})\in R_{n-3}$. Since $u_{m}$ is linear in $\xi_{n-2}$ and divides $\mathcal{P}^{i}(u_{m})\in R_{n-3}$, we conclude that $\mathcal{P}^{i}(u_{m})=0$ for $0<i<q^{m-1}$.

### v3

@@ 6. The Invariants of $\text{O}^{+}_{{4}}(\mathbb{F}_{q})$}}
From (c) we have
$u_{m}=(\xi_{n-1}+c_{n-1,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-1-i}+c_{n-1-i,m})(u_{m-1}d_{i,m-1})^{q}$
with $c_{j,m}\in R_{j-1}$ and $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n}\in R_{n-1}$. Therefore
$F:=\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-i}+c_{n-i,m})(u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n})\in R_{n-1}.$
The coefficient of $\xi_{n}$ in the expression defining $F$ is
$\sum_{i=1}^{m}(-1)^{i}(\xi_{n-i}+c_{n-i,m})(u_{m-1}d_{i-1,m-1})^{q}=-u_{m}.$
Thus
$F=-u_{m}\xi_{n}+\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-i}+c_{n-i,m})u_{m}d_{i,m}\right)$
and
$\xi_{n}=-F/u_{m}+\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-i}+c_{n-i,m})d_{i,m}\right).$
Since $F$ is divisible by $u_{m}$ in $S_{m}^{G_{m}}$ and $F\in R_{n-1}$, we have $F\in\ker(\Phi_{n-1,m-1})$ (see Lemma 4.2 and Remark 4.7). From part (a), $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$. Therefore $F$ is divisible by $u_{m}$ in $R_{n-1}$. Define $c_{n,m}:=F/u_{m}\in R_{n-1}$. This gives
$\xi_{n}=-c_{n,m}+\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-i}+c_{n-i,m})d_{i,m}\right)$
as required. ∎
7.2. Proof of Parts (a), (b) and (c)
The proof is by induction on $m$. The base cases $m=2$ is given in Section 6. By the induction hypothesis $S_{m-1}^{G_{m-1}}$ is generated by $\{\xi_{1},\ldots,\xi_{n-3},d_{1,m-1},\ldots,d_{m-1,m-1}\}$. The Hook subgroup for $G_{m}$ is a normal subgroup of the pointwise stabiliser $(G_{m})_{x_{m}}$. If we identify $G_{m-1}$ with the pointwise stabiliser of $\{x_{m},y_{m}\}$ in $G_{m}$, then $(G_{m})_{x_{m}}=G_{m-1}H$. Arguing as in the proof of Theorem 5.11 and using Theorem 5.5, we see that $S_{m}^{(G_{m})_{x_{m}}}$ is generated by
$\{N(y_{m}),x_{m},\xi_{1},\ldots,\xi_{n-2},\psi_{[m,1]}(\sigma(d_{1,m-1})),\ldots\psi_{[m,1]}(\sigma(d_{m-1,m-1}))\}.$
In this context, the compliance relation (Lemma 5.3) allows us to rewrite $\xi_{n-1}$ in terms of the generators of $S_{m}^{(G_{m})_{x_{m}}}$. For degree reasons, the resulting expression is linear in the $\psi_{[m,1]}(\sigma(d_{j,m-1}))$, giving
$\xi_{n-1}=x_{m}N(y_{m})+\lambda_{0}+\sum_{j=1}^{m-1}\lambda_{j}\psi_{[m,1]}(\sigma(d_{j,m-1}))$
with $\lambda_{j}\in R_{n-2}[x_{m}]$. Since $u_{m}=x_{m}N(y_{m})\psi_{[m,1]}(\sigma(u_{m-1}))$, we have
$u_{m}=\psi_{[m,1]}(\sigma(u_{m-1}))\left(\xi_{n-1}-\lambda_{0}-\sum_{j=1}^{m-1}\lambda_{j}\psi_{[m,1]}(\sigma(d_{j,m-1}))\right).$
We know that $\psi_{[m,1]}$ and $\sigma$ are algebra maps. Furthermore, $\psi_{[m,1]}(\sigma(f))=\psi_{[m,1]}(f)$ for $f\in R_{n-2}$. By induction $u_{m-1}$ and $u_{m-1}d_{j,m-1}$lie in $R_{n-2}$. Therefore
$u_{m}=(\xi_{n-1}-\lambda_{0})\psi_{[m,1]}(u_{m-1})-\sum_{j=1}^{m-1}\lambda_{j}\psi_{[m,1]}(u_{m-1}d_{j,m-1}).$
Write $\lambda_{j}=\overline{\lambda_{j}}+x_{m}\lambda_{j}^{\prime}$ with $\overline{\lambda_{j}}\in R_{n-2}$ and $\lambda_{j}^{\prime}\in R_{n-2}[x_{m}]$ and observe that $\psi_{[m,1]}(\lambda_{j})=\psi_{[m,1]}(\overline{\lambda_{j}})$. Using Lemma 4.2, $\psi_{[m,1]}(\psi_{[m,1]}(f))=\psi_{[m,1]}(f)^{q}$ for $f\in R_{n-2}$. Applying $\psi_{[m,1]}$ to the above expression and using the fact that $\psi_{[m,1]}(u_{m})=0$ gives
$0=\psi_{[m,1]}(\xi_{n-1}-\overline{\lambda_{0}})\psi_{[m,1]}(u_{m-1})^{q}-\sum_{j=1}^{m-1}\psi_{[m,1]}(\overline{\lambda_{j}})\psi_{[m,1]}(u_{m-1}d_{j,m-1})^{q}.$
@@ 7.2. Proof of Parts (a), (b) and (c)
(7) $\widetilde{u}_{m}:=(\xi_{n-1}-\overline{\lambda_{0}})u_{m-1}^{q}-\sum_{j=1}^{m-1}\overline{\lambda_{j}}(u_{m-1}d_{j,m-1})^{q}\in R_{n-1}\cap\ker(\psi_{[m,1]}).$
Using Lemma 4.2, $u_{m}$ divides every element of $R_{n-1}\cap\ker(\psi_{[m,1]})$. Therefore $u_{m}$ divides $\widetilde{u}_{m}$. Since ${S\mbox{-}\deg}(u_{m})={S\mbox{-}\deg}(\widetilde{u}_{m})$, we see that $\widetilde{u}_{m}$ is a scalar multiple of $u_{m}$. As an element of $R_{n-1}$, $\widetilde{u}_{m}$ is linear in $\xi_{n-1}$ with coefficient $u_{m-1}^{q}\not=0$. Therefore $\widetilde{u}_{m}$ is a non-zero scalar multiple of $u_{m}$. From this we conclude that $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$. We claim that that $\widetilde{u}_{m}=u_{m}$. To see this, first observe that $\operatorname{LT}(u_{m})=\prod_{i=1}^{m}y_{i}^{q^{m+i-2}}x_{i}^{q^{m-i}}=:\beta$ using either the lex or grevlex order on $S_{m}$. We will show that $\beta$ appears in $\widetilde{u}_{m}$ with coefficient one. The factor $y_{m}^{q^{n-2}}x_{m}$ can only come from a factor of $\xi_{n-1}$. Therefore, using that the expression for $\widetilde{u}_{m}$ as an element of $R_{n-1}$, we see that $\beta$ can only appear as a term in $\xi_{n-1}u_{m-1}^{q}$. The factor $y_{m-1}^{n-3}x_{m-1}^{q}$ necessarily comes from a factor of $\xi_{n-3}^{q}$ appearing in a term in $u_{m-1}^{q}$. Iterating this process and using induction, we see that $\xi_{n-1}\xi_{n-3}^{q}\cdots\xi_{1}^{q^{m-1}}$ appears with coefficient one in $\widetilde{u}_{m}\in R_{n-1}$ and $\beta$ appears with coefficient one in $\widetilde{u}_{m}\in S_{m}$. Therefore $\widetilde{u}_{m}=u_{m}$. Applying $\mathcal{P}^{e(i,m)}$ to this expression shows that $u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(u_{m})\in R_{n}$ and $u_{m}d_{i,m}-\xi_{n}\mathcal{P}^{e(i,m)-q^{n-2}}(u_{m-1}^{q})\in R_{n-1}$. Since $e(i,m)-q^{n-2}=qe(i-1,m-1)$, we have $u_{m}d_{i,m}-\xi_{n}(u_{m-1}d_{i-1,m-1})^{q}\in R_{n-1}$.
@@
As an element of $S_{m}$, $u_{m}$ is a product of pairwise relatively prime linear forms. Therefore $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $S_{m}$. If $\mathcal{P}^{i}(u_{m})\in R_{n-1}$, then
$\mathcal{P}^{i}(u_{m})\in\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$
and $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $R_{n-1}$(see Remark 4.7). For $i<q^{n-2}$, $\mathcal{P}^{i}(u_{m})\in R_{n-1}$ and, therefore, $\mathcal{P}^{i}(u_{m})/u_{m}\in R_{n-1}$ . By induction, $\mathcal{P}^{i}(u_{m-1})=0$ for $0<i<q^{m-2}$. Therefore, for $0<i<q^{m-1}$, $\mathcal{P}^{i}(u_{m-1}^{q})=0$ and $\mathcal{P}^{i}(u_{m})\in R_{n-2}$. Since $u_{m}$ is linear in $\xi_{n-1}$ and divides $\mathcal{P}^{i}(u_{m})\in R_{n-2}$, we conclude that $\mathcal{P}^{i}(u_{m})=0$ for $0<i<q^{m-1}$.